# HD-OOD3D: Supervised and Unsupervised Out-of-Distribution object detection in LiDAR data

### v2

@@ IV-B1 Evaluation set
Using the ”Open” Scans subset, which includes only scans with unknown objects, significantly improved the AUPR-E metric from 12.6 to 20.0 for HD-OOD3D. Due to the known/unknown class imbalance, using all scans adds many more potential false positives without any unknown ground truth, hindering all methods across the board. Consequently, our final evaluation used the ”open” subset—only the evaluation scans containing at least one unknown object—to better assess performance. It is important to note that, as shown in Table IV, the chosen hyperparameters do not maximize the performance of HD-OOD3D in isolation, but rather provide a balanced and fair benchmark for all methods.
@@ IV-B2 Distance and Threshold Settings for OOD Detection
The choices of distance metric $d_{\text{thresh}}$ and detection module threshold $\delta_{\text{thresh}}$ play a critical role in OOD Detection.
Adjusting the distance threshold for matching predicted boxes to ground truth objects influences the inclusion of predictions in metric calculations. Specifically, when using a larger $d_{\text{thresh}}$ (e.g., 2.0m instead of 0.5m), more predictions, mainly those correctly identifying unknown classes, are included. This increase in detection hits—predicted boxes within the range of ground truth—allows for a more comprehensive evaluation of the model’s ability to detect unknown objects. However, this also introduces a risk of misassigning predictions to known objects, potentially lowering AUROC and FPR-95 while improving AUPR-E, as this metric directly reflects the precision in detecting unknowns.
@@
$\delta_{\text{thresh}}$ $d_{\text{thresh}}$ Hits (U)Hits (C)AUROC $\uparrow$ FPR-95 $\downarrow$ AUPR-E $\uparrow$ AUPR-S $\uparrow$
@@
TABLE IV: Impact of thresholds $\delta_{\text{thresh}}$ and $d_{\text{thresh}}$ on Detection Hits (for unknown $U$ instances and closed known $C$ instances), and OOD Metrics for Open and Closed Ground Truth Objects in the Argoverse2 Dataset. Underlined parameters represent our final hyperparameter choices.
@@ IV-B3 Sorting method
V Methods for Generating Additional Unknown data for training the MLP
@@ V-A Different Representations of Out-of-Distribution Objects
(b) External Data
@@
Figure 2: Unknown generation methods for training two-stage OOD detection methods
In contrast to the single-generation approach used in MMOOD3D [16], we propose several alternative methods to synthesize out-of-distribution (OOD) objects, specifically a Top-K labelling approach. We compare these different representations of out-of-distribution objects for training, which we create through several means:
  • Gaussian noise [9]: we perturb the extracted embedding of ID objects with Gaussian noise, adding a normal distribution $\mathcal{N}(0,1)$ to the extracted embeddings.
@@
  • Resizing: create out-of-distribution objects by resizing in-distribution (ID) known objects to modify their scales along all three axes, similarly to [16]. Each known object with at least 5 points has a 50% probability of being resized, where for each axis, we randomly apply a scaling factor between $[0.1,0.5]$ and $[1.5,3.0]$.
  • Our 3D Top-K autolabelling : inpired by [3] for open 2d objet detection in images, we propose in 3D to use the top-K most confident predictions from the base detector, which don’t overlap with ground truth known objects, as OOD objects. In this approach, we keep the scans from the datasets containing real OOD distribution, as we expect these to become autolabelled. Since other methods do not rely on real OOD data for training the MLP, we remove those LiDAR scans for other methods.
We illustrate these different generation methods in Figure 2 and compare these various representations of OOD objects and their impact on the training and performance of our additional module on real validation OOD data.
@@ V-B Results
The Resizing method obtained an AUROC of 77.2, with a relatively low AUPR-E of 6.3. Compared to the original paper, these low results highlight how this simple approach might not be suitable for the more diverse Argoverse 2. The Gaussian Noise method performed poorly across all metrics, with an AUROC of 66.5 and a very high FPR-95 of 84.5, making it only slightly better than random chance.
Unsupervised Unknowns AUROC $\uparrow$ FPR-95 $\downarrow$ AUPR-E $\uparrow$ AUPR-S $\uparrow$
@@
TABLE V: Comparison of AUROC, FPR-95, AUPR-S, AUPR-E and AP Metrics for different unknown generation methods on Argoverse 2. methods represent adaptations of data augmentation methods.
ModelNet Injection, which incorporates synthetic objects from the ModelNet dataset as OOD data, showed better performance with an AUROC of 70.6 and an AUPR-E of 5.0. We believe the difficulties in OOD detection for this method stem from the synthetic nature of the dataset, which suggests that while this method might be interesting [14], one has to be careful about the quality of these instances and the domain differences between these and the real unknowns. PointMixUp struggled with a high FPR-95 of 77.6 and a low AUPR-E of 4.6, indicating challenges in detecting unknown objects despite the novel hybrid representations. The Top-K Labelling (K=5) method, which labels the top 5 most confident predictions not overlapping with ground truth known objects as OOD, emerged as a notable performer. It achieved an AUROC of 81.9 and the highest AUPR-S of 99.6, demonstrating good identification of known objects. Despite its higher FPR-95 of 68.2, Top-K Labelling showed promise in balancing the detection of known and unknown objects by leveraging real-world OOD distributions in the dataset. We experimented with different K values (K= $\{2,5,10,20\}$) and found that K=5, close to the average number of unknown class instances per frame, performed best across all metrics. Overall, the remaining gap between the two-stage oracle MLP and this highest-performing unknown generation method suggests that future works should focus on improving this pseudo-labelling process.
@@ VI CONCLUSIONS
In this work, we presented a new two-stage method for detecting unknown objects called HD-OOD3D. By exploiting a sufficiently large dataset such as Argoverse 2 to train the unknowns in a supervised manner, we demonstrated the superiority of two-stage methods over single-stage methods, particularly our proposed HD-OOD3D, which utilizes more input representations for better object categorization. We then broke down the classic evaluation protocol for unknown objects to highlight the importance of certain hyperparameter choices for a fair assessment of methods. Finally, scaling the learning of unknown objects can only be achieved through unsupervised training. We compared different approaches for generating unknowns by using only the annotations of known objects from Argoverse 2. We showed that top-5 auto-labelling offers more promising performance perspectives compared to resizing. Ultimately, the task of detecting unknown objects remains a challenging one, with current performance far from what would enable deployment on robotics platforms, making it a field that still requires further research to reach maturity.

### v3

@@ IV-B1 Evaluation set
Using the ”Open” Scans subset, which includes only scans with unknown objects, significantly improved the AUPR-E metric from 12.6 to 20.0 for HD-OOD3D. Due to the known/unknown class imbalance, using all scans adds many more potential false positives without any unknown ground truth, hindering all methods across the board. Consequently, our final evaluation used the ”open” subset, only the evaluation scans containing at least one unknown object, to better assess performance. It is important to note that, as shown in Table IV, the chosen hyperparameters do not maximize the performance of HD-OOD3D in isolation, but rather provide a balanced and fair benchmark for all methods.
@@ IV-B2 Distance and Threshold Settings for OOD Detection
The choices of detection module threshold $\delta_{\text{thresh}}$ (threshold above which the boxes are considered as objects by the base detector) and distance metric $d_{\text{thresh}}$ (distance used for associating predicted boxes with ground-truth boxes) play a critical role in OOD Detection.
Adjusting the distance threshold for matching predicted boxes to ground truth objects influences the inclusion of predictions in metric calculations. Specifically, when using a larger $d_{\text{thresh}}$ (e.g., 2.0m instead of 0.5m), more predictions, mainly those correctly identifying unknown classes, are included. This increase in detection hits, predicted boxes within the range of ground truth, allows for a more comprehensive evaluation of the model’s ability to detect unknown objects. However, this also introduces a risk of misassigning predictions to known objects, potentially lowering AUROC and FPR-95 while improving AUPR-E, as this metric directly reflects the precision in detecting unknowns.
@@
$\delta_{\text{thresh}}$ $d_{\text{thresh}}$ Hits (U)% Hits (C)% AUROC $\uparrow$ FPR-95 $\downarrow$ AUPR-E $\uparrow$ AUPR-S $\uparrow$
@@
TABLE IV: Impact of thresholds $\delta_{\text{thresh}}$ and $d_{\text{thresh}}$ on Detection Hits (for unknown $U$ instances and closed known $C$ instances), and OOD Metrics for Unknown and Known Ground Truth Objects in the Argoverse 2 dataset. Underlined parameters represent our final hyperparameter choices.
@@ IV-B3 Sorting method
V Comparison of Methods for Unsupervised Learning of Unknown 3D Objects
@@ V-A Different Representations of Out-of-Distribution Objects
(b) External Data from ModelNet [23]
@@
Figure 2: Generation Methods for Unknown Objects in 3D LiDAR to Supervise Two-Stage OOD Detection
In contrast to the single-generation approach used in MMOOD3D [17], we propose several alternative methods to synthesize out-of-distribution (OOD) objects. We compare these different representations of out-of-distribution objects for training, which we create through several means:
  • Gaussian noise [9]: we perturb the extracted embedding of ID objects with Gaussian noise, adding a normal distribution $\mathcal{N}(0,1)$ to the extracted embeddings. This approach allows us to inject variability into the feature space, thereby creating synthetic representations of unknown objects.
@@
  • Resizing: we create out-of-distribution objects by resizing in-distribution (ID) known objects to modify their scales along all three axes, similarly to [17]. Each known object with at least 5 points has a 50% probability of being resized, where for each axis, we randomly apply a scaling factor between $[0.1,0.5]$ and $[1.5,3.0]$. This simple transformation creates new, out-of-distribution representations by altering the spatial dimensions of familiar objects.
  • Our 3D Top-K autolabelling : inpired by [3] for open 2d objet detection in images, we propose in 3D to use the top-K most confident predictions from the base detector, which don’t overlap with ground truth known objects, as OOD objects. In this approach, we keep the scans from the datasets containing real OOD distribution, as we expect these to become autolabelled.
We illustrate these different generation methods in Figure 2 and compare these various representations of OOD objects and their impact on the training and performance of our additional module on real validation OOD data. All trainings are done with our approach HD-OOD3D of Section III-A.
@@ V-B Results
The Resizing method obtained an AUROC of 77.2, with a relatively low AUPR-E of 6.3. These low results highlight how this simple approach might not be suitable for the more diverse Argoverse 2. The Gaussian Noise method performed poorly across all metrics, with an AUROC of 66.5 and a very high FPR-95 of 84.5, making it only slightly better than random chance.
Methods AUROC $\uparrow$ FPR-95 $\downarrow$ AUPR-E $\uparrow$ AUPR-S $\uparrow$
@@
TABLE V: Comparison of AUROC, FPR-95, AUPR-E, and AUPR-S metrics for different unknown generation methods on Argoverse 2. methods represent adaptations of data augmentation methods. The yellow cells indicate methods with an AUPR-S score below 99, which means that the base detector has lost performance on the known classes. GT Unknowns uses the labels of the unknown objects to supervise the MLP and can be regarded as an oracle.
ModelNet Injection, which incorporates synthetic objects from the ModelNet dataset as OOD data, showed better performance with an AUROC of 70.6 and an AUPR-E of 5.0. We believe the difficulties in OOD detection for this method stem from the synthetic nature of the dataset, which suggests that while this method might be interesting [15], one has to be careful about the quality of these instances and the domain differences between these and the real unknowns. PointMixUp struggled with a lower AUROC of 72.1 compared to the Resizing method with 77.2, even though it remains the best method for the FPR-95 and AUPR-E metrics compared to previous methods.
The Top-K Labelling method, which labels the top K most confident predictions not overlapping with ground truth known objects as OOD, emerged as a notable performer. It achieved an AUROC of 81.9 and the highest AUPR-S of 99.6, demonstrating good identification of known objects. Despite its higher FPR-95 of 68.2, Top-K Labelling showed promise in balancing the detection of known and unknown objects by leveraging real-world OOD distributions in the dataset. We experimented with different K values (K= $\{2,5,10,20\}$) and found that K=5 (reported in Table V), close to the average number of unknown class instances per frame, performed best across all metrics. Overall, the remaining gap between the two-stage oracle MLP (GT Unknowns in Table V) and this highest-performing unknown generation method suggests that future works should focus on improving this pseudo-labelling process.
@@ VI CONCLUSIONS
In this work, we presented a new two-stage method for detecting unknown objects called HD-OOD3D. By exploiting a sufficiently large dataset such as Argoverse 2 to train the unknowns in a supervised manner, we demonstrated the superiority of two-stage methods over single-stage methods, particularly our proposed HD-OOD3D, which utilizes more input representations for better object categorization. We then broke down the classic evaluation protocol for unknown objects to highlight the importance of certain hyperparameter choices for a fair assessment of methods. Finally, scaling the learning of unknown objects can only be achieved through unsupervised training. We compared different approaches for generating unknowns by using only the annotations of known objects from Argoverse 2. We showed that top-K auto-labelling offers more promising performance perspectives compared to resizing. Ultimately, the task of detecting unknown objects remains a challenging one, with current performance far from what would enable deployment on robotics platforms, making it a field that still requires further research to reach maturity.